OpenStudy (anonymous):

If a<=b<=c<=d<=e<=110 and the average of a,b,c,d,e is 100 what is the least possible value of a? options: 0 , 20 , 40, 60, 80

HanAkoSolo (jamierox4ev3r):

the layout of this is pretty confusing, is it less than or equal to?

OpenStudy (reemii):

the question to start is: if a=0, is it possible to have an average of 100 ? (test: b=c=d=e=110, average: 440/5 < 100. oops no!) if a=20, is it possible to have an average of 100? (test: b=c=d=e=110, average (440+20)/5 .. etc)

HanAkoSolo (jamierox4ev3r):

Basically, is what you wanted to portray this: \[\ge or\]

OpenStudy (anonymous):

@reemii thanks that makes sense!

OpenStudy (reemii):

I think the answer is 60. the sum can be at most 500, divided by 5 you have exactly 100.

OpenStudy (phi):

a<=b<=c<=d<=e<=110 for a to be smallest, you want b,c,d,e to be largest. the largest each can be is 110, the average in that case would be (a+110+110+110+110)/5 or (a+440)/5 if this equals 100 you have (a+440)/5 = 100 solve for a

OpenStudy (anonymous):

@phi thanks that is a great answer

OpenStudy (reemii):

@arykk it's the better one for sure :)